A large object of mass m 10 kg is initially at rest. It splits into two unequal pieces. The smaller piece has a speed of 5 m/s to the right and the larger piece has a speed of 2 m/s to the left. What is the mass of the smaller piece? (a) 0.25 kg (b) 2.5 kg (c) 2.9 kg (d) 4 kg (e) You need to know one of the masses to solve for the other
Answer #1

m1 + m2 = 10

By the conervation of momentum we have

mV = m1u1 + m2u2

10*0 = m1*5 + (10-m1)*(-2)

m1 = 20/7 = 2.857 = 2.9 kg

so option c) 2.9 kg is the correct answer